If x+1/x-3 =0 then find the value of
(1)X^2+1/x^2
(2)x^4+1/x^4
(3)x^5+1/x^5

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
We can write it as
 
x + 1/x = 3
 
Now square
 
x² + 1/x² + 2 = 9
 
Hence x² + 1/x² = 7
 
Similarly you can find out rest parts